OpenStudy (anonymous):

Factor 35d + 5d2 - 30. 5(d2- 7d + 6) 5(d2- 7d - 6)****** 5(d2+ 7d - 6) I think is right *****

OpenStudy (anonymous):

You are unfortunately incorrect. Notice the +/- signs!

OpenStudy (anonymous):

so the 2nd one

OpenStudy (anonymous):

\[5d^2+35d-30=5(\frac{5d^2}{5}+\frac{35d}{5}-\frac{30}{5})\]

OpenStudy (anonymous):

so I was rong

OpenStudy (anonymous):

Yes. But just slightly wrong. You didn't notice that 35d is positive

OpenStudy (anonymous):

okay